I dont understand how to solve for x in this equation.
Problem :
125^10x-9 = (1/625)^2x-5
Could someone please help... Thanks

Answer by Alan3354(69443)

125^10x-9 = (1/625)^2x-5
--------------
125 = 5^3
1/625 = 5^-4
5^(3*(10x-9)) = 5^(-4*(2x-5))
Since the base is the same, the exponents are equal.
3(10x-9) = -4(2x-5)
You can take it from there.